OpenStudy (anonymous):

ok so this is the problem (-4,-19) and (-15,-20) which come to y=1/11 +76/11 so tell me how i begin to graph it in the first place please

OpenStudy (amistre64):

what was wrong with the other posting?

OpenStudy (anonymous):

i didnt see the other posting

OpenStudy (amistre64):

do you want to graph this according to the equation given or the points given?

OpenStudy (anonymous):

so can you help me again

OpenStudy (amistre64):

i can try :)

OpenStudy (anonymous):

well it says

OpenStudy (anonymous):

to find the slope and graph it

OpenStudy (anonymous):

so the slope would be 77\11 correct?

OpenStudy (amistre64):

let me chk your slop then (-4,-19) -(-15,-20) ----------- 11 , 1 ; slope = y/x = 1/11, that parts good the equation will be, in slope intercept form: y = (1/11)x --4/11 -19(11)/11 = (1/11)x + (4-209)/11 = (1/11)x - 205/11 as i see it

OpenStudy (amistre64):

these are some rather difficult numbers to play with; so we can use one of the points given and step it thru with the slope

OpenStudy (amistre64):

id start at: (-4,-19) and move, according to the slope, up 1 and right 11 to get to; (-4, -19) +1, +11 --------- (-3, -8)

OpenStudy (amistre64):

i got me x and y backwards .... it should be (-4, -19) +11, +1 --------- ( 7, -18) as a new point

OpenStudy (amistre64):

something akin to this picture ...

OpenStudy (anonymous):

so y do we add the second set of numbers to the first answer?

OpenStudy (amistre64):

the second set of numbers is how the slope directs us to move from one point to another in order to find another point on a slope of 1/11 we have to move up 1 and right 11

OpenStudy (amistre64):

it is simply following the directions given by the slope: 1 movement along the y -- = -------------------- 11 movement along the x

OpenStudy (anonymous):

thanks